# Generating Temporal Networks with the Ascona Model

### v1

@@ Abstract
We introduce a new sampling method for continuous-time temporal networks based on queueing processes. In particular, we focus on a Markovian version of the model where the links between nodes are Poisson distributed in time and have exponential duration. We highlight the stochastic properties of these temporal structures and leverage them to design synthetic temporal networks with a controllable level of smoothness, which follow patterns relevant for the validation and interpretation of methods for community, scale, change-point, and periodicity detection. Additionally, we show that imposing assortativity constraints on the samples leads to a continuous-time generalization of stochastic block models. Finally, we describe how variations of the model can be used to sample alternative types of structure and temporal networks, especially discrete-time ones.
@@ 1 Introduction
Temporal (or dynamic) networks/graphs are extensions of graphs that incorporate temporal information. Various mathematical objects can be considered to be temporal networks, and different formalisms have been proposed to describe them (e.g., [casteigts_time-varying_2012, holme_temporal_2012, latapy_stream_2018]). These objects can be classified on a spectrum of increasing complexity based on the granularity of the encoded temporal information [rossetti_social_2015, rossetti_community_2018]. On one side, we have static graphs, which aggregate in one structure all temporal information; on the other side, there are stream graphs and link streams that avoid any aggregation. Somewhere in between the two sides lie snapshot networks and discrete-time stream graphs/link streams.
@@
Temporal networks are relevant in modeling many complex systems [masuda_guide_2016], and, as for their static counterparts, understanding their communities/clusters is a fundamental problem in network science. The community structure evolution is typically characterized by archetypal structural changes [cazabet_challenges_2023], called events, or combinations thereof [failla_describing_2024]. Therefore, sampling methods that allow to model these archetypes are of paramount importance for the generation of quality synthetic data with controlled ground-truth. One needs these synthetic networks, for instance, for the validation and interpretation of algorithms tasked for community, scale, change-point, and periodicity detection. A review of the main existing methods, particularly relevant for community detection benchmarks, can be found in [cazabet_evaluating_2021].
@@
For some applications, one would also want to sample networks that are less artificial and instead are closer to real-world datasets. Recently proposed techniques in this direction (e.g., [gauvin_randomized_2022, longa_efficient_2022, presigny_building_2021]) aim at producing surrogate temporal graphs, structures that maintain some or most properties of a reference (often real-world) temporal network. However, these methods do not provide sufficient control to design a controlled experimental setup, as they bypass the explicit formulation of a modeling mechanism by relying on a pre-existing structure.
@@
Agent-based models (e.g., [perra_activity_2012, snijders_introduction_2010]) are an alternative option. They leverage control of microscopic dynamics of a complex system and combine them so that, in the aggregate, complex phenomena emerge. However, often, the understanding of these models relies heavily on simulations, as the underlying dynamics are complex, leading to mathematical intractability, hence limited stochastic and statistical control.
@@
Most of the temporal network literature, especially on sampling methods, has focused on snapshot networks or discrete-time structures, leaving out the more complex continuous-time ones (a notable exception being [bovet_flow_2022]). The main contribution of this article is the introduction of a sampling method for continuous-time temporal networks, and the explanation of how to use it to sample data reflecting the above-mentioned event patterns, plus a continuous-time generalization of the stochastic block model (SBM) [abbe_community_2018].
@@
The model works in two fundamental steps: in the first, we sample when nodes are connected with a link, and in the second, we sample which nodes to assign to each link sampled in the previous step. Separating the temporal from the connectivity aspect offers benefits in terms of implementation and clarity, while not sacrificing expressivity. This separation mechanism also helps to structure the mathematical analysis of a temporal network (model), as the two aspects can be considered orthogonal to each other, hence treated separately, as in [barucca_maximum_2025]. Connectivity can exhibit almost arbitrary levels of complexity and is responsible for various structural aspects such as assortativity, disassortativity, and nestedness. A complete analysis of its properties from a stochastic and statistical perspective is therefore challenging. The temporal aspect appears to be simpler; however, surprisingly, its probabilistic and statistical analysis has been neglected. This work addresses this gap: its key idea is to describe the appearance and disappearance of nodes’ connections in terms of queueing processes, and then focus on the evolution of the total number of active connections as a function of time to derive probabilistic and statistical results.
@@
Queueing processes were originally introduced to describe the flow of customers (jobs, particles, calls) arriving at a service facility. It is possible to approximate the evolution of most queueing systems with Markov chains. The advantage of these is that it is possible to calculate the stationary distribution of a chain by solving a linear system of equations called the balance conditions [walrand_introduction_1988].
@@
The structure of this work is as follows. In Section˜2 we introduce the notation and describe the objects studied in the article. It is divided into two parts: the first is devoted to formalisms of temporal networks, the second to queueing processes. In Section˜3 we introduce the Ascona model, a temporal networks model based on queueing theory, and then use a Markovian parametrization of it to generate link streams with desired temporal evolutions. Based on this Markovian framework, we describe a link stream generalization of SBMs, and compare it to their currently most popular temporal extensions, dynamic SBMs. We conclude the section with a description of variations of the Ascona model. The end of the article, Section˜4, contains a discussion of the presented results and highlights possible future research directions stemming from this work.
@@ 2 Theoretical Framework
In this section, we describe the notation and objects we need to introduce the Ascona model. In the first part, we describe various formalisms of temporal networks following the framework presented in [latapy_stream_2018], and in the second part, general queueing processes, focusing on the $M/M/\infty$ queue, also known as the ample-server model [gross_fundamentals_2008, Ch. 2.7].
@@ 2.1.1 Link Streams
We say that a stream $S^{\prime}=(T^{\prime},V^{\prime},E^{\prime})$ is a sub-stream of $S=(T,V,E)$ if, intuitively, $T^{\prime}\subset T,V^{\prime}\subset V,E^{\prime}\subset E$.
@@ 2.1.2 Snapshot Networks
The length of each slice is determined based on a compromise: one needs slices large enough to ensure that each snapshot captures significant structural information, but large slices lead to losses of temporal information, since all interactions within the same slice are merged. In addition, several or even varying slice durations may be relevant [latapy_stream_2018].
@@ 2.2 The $M/M/\infty$ Queue
Kendall’s notation, originating from [kendall_stochastic_1953], concisely encodes these features as
@@
We are mainly interested in one specific queue, the $M/M/\infty$ one. The ‘$M$’ is a standard symbol in Kendall notation indicating Markovian processes; therefore, the model can be interpreted as a birth (or immigration)-death process where arrivals form a Poisson process with rate $\lambda>0$, and deaths are i.i.d. exponential with rate $\mu>0$ [asmussen_applied_2003]. There are countably infinitely many servers, hence no waiting occurs: each arrival immediately starts the death process.
@@
$p_{m}^{\prime}(t)$ $=\lambda p_{m-1}(t)-(\lambda+n\mu)p_{m}(t)+(n+1)\mu p_{m+1}(t),\qquad m\geqslant 1.$ (2.11)
@@ 2.2.1 Process Evolution
It is known (e.g., [robert_stochastic_2013, Ch. 6]) that, assuming that $M(0)=0$, $M(t)$ is Poisson with a time–dependent mean
@@
If we stop sampling elements after a certain time $t_{e}$, then the process is truncated, and the mean follows the equation
$m(t)=\begin{cases}0,&t<t_{s},\\[4.0pt] \lambda\mu\left(1-e^{-(t-t_{s})/\mu}\right),&t_{s}\leq t\leq t_{e},\\[4.0pt] \lambda\mu\left(e^{-(t-t_{e})/\mu}-e^{-(t-t_{s})/\mu}\right),&t>t_{e}.\end{cases}$ (2.18)
@@ 2.2.2 Stationary Regime
One could also recover $\mathbb{E}[M(\infty)]=\frac{\lambda}{\mu}$ directly from Little’s law (see [little_proof_1961] and [walrand_introduction_1988, Thm. 3.4.1]).
We now derive an estimate for the time $t^{\star}$ at which we can assume that $m(t)$ is close enough to the stationary value. We use this estimate later in Section˜3.4 and Section˜3.5.
@@ 2.2.3 Head of the Queue
We consider an $M/M/\infty$ queue that starts empty at time $t=0$, with arrival rate $\lambda>0$ (Poisson arrivals) and i.i.d. service times $S\sim\mathrm{Exp}(\mu)$ of mean $\mu>0$.
@@
In the case of $N$ queue samples, the standard deviation of the sample mean is $\sqrt{\frac{m(t)}{N}}$, so it takes more time to establish closeness to stationarity. The computation is equivalent to that of the single sample case. This heuristic argument can be made rigorous in terms of the total variation distance, but a more refined estimation based on mixing times [levin_markov_2017] is possible, allowing one to control closeness to stationarity to an arbitrary level of precision.
@@ 2.2.4 Number of Elements in the Queue
The number of elements that join the queue in a time interval $t_{1},t_{2}]$ is, by design, Poisson distributed with average $\lambda(t_{2}-t_{1})$. The number of elements in the queue in the same interval is slightly different. In addition to new elements joining in that interval, we need to count how many started before and ended after $t_{1}$. This amounts approximately to the sum of two independent random variables, one Poisson with average $\lambda(t_{2}-t_{1})$ and the other Poisson with average $m(t_{1})$, which by superposition is again Poisson with average $\lambda(t_{2}-t_{1})+m(t_{1})$. In the stationary regime, this becomes a Poisson random variable with average $\lambda(t_{2}-t_{1})+\lambda/\mu$. We use these facts later in Section˜3.5.
3 The Ascona Model and EDEDE Structures
We design a link streams model based on queueing processes. In the absence of a previously established name, we will refer to it as the Ascona model/method for clarity of the presentation and convenience.
@@ 3.1 The Ascona Model
The method relies on the sampling of fundamental sub-stream units called queue blocks sharing the same vertex set $V$ and time domain $T$. Each block links set $E$ is determined in two separate steps: in the first, we create a timeline of the links belonging to it, and in the second, we assign each link to a pair of nodes. The global network is then obtained by combining queue blocks sequentially and in parallel. We first start by describing the creation of a single queue block.
3.1.1 Temporal Structure
The number of links in a queue block, their start and ending times are sampled based on a queueing process $A/S/c/K/\bar{K}/D$ in the interval $0,t_{e}]$, plus the tail obtained by stopping the admission of new elements in the queue after $t_{e}$. The choice of the queue parameters can be based on ad hoc properties of the repartition in time of links, or prior information about it observed in a reference network. The same goes for the choice of which section of the queue to retain (e.g., keep only the head, the tail, or the stationary part of the queue as discussed later in Section˜3.4) and where to situate the block in the time domain (i.e., applying a translation factor to all elements of the queue). In this work, we deal exclusively with ad hoc selection and placements of queue blocks; however, one could imagine randomizing this process in various ways.
The maximal possible number of links at each time is $\binom{N}{2}$, however we can still work with queues with $K>\binom{N}{2}$ if we have a mechanism to stop sampling links when the queue is congested, as we explain later. For instance, in Section˜3.2, we discuss a fundamental queue parametrization, which assumes Markovian arrival and service times, plus infinitely many servers.
3.1.2 Connectivity Structure
After sampling the number of links and their starting and ending times, we need to assign each link to a pair of nodes. We do it stochastically by sampling according to a probability distribution on $V\times V$. We can encode such a distribution in an $N\times N$ matrix whose entries add up to one. We will refer to it as the connectivity probability matrix.
Alternatively, it is often convenient to split the selection process into two. In the first step, a node is picked up randomly according to a probability distribution whose support is a subset of $V$. For instance, a convenient choice would be the uniform distribution, which has maximal entropy and therefore can be interpreted as uninformative from a Bayesian perspective. One could bias this choice based on prior knowledge of (potentially uneven) nodes’ activity, a crucial point in [perra_activity_2012]. In the second step, we sample the second node based on a probability distribution on $V$ dependent on the first node. We can encode these distributions in an $N\times N$ row-stochastic matrix. We will refer to it as the connectivity conditional probability matrix, as the $(i,j)$-th entry corresponds to the probability of selecting $v_{j}$ for a link, given that the first sampled node was $v_{i}$.
In both cases, we impose that the diagonal elements of the connectivity matrices are all equal to zero to avoid self-links.
We show a concrete example of two dynamic networks sampled with the Ascona method in Figure˜2. The two samples share the same temporal structure but have different connectivity. The shared temporal structure is sampled based on an $M/M/\infty$ queue, whose construction is represented in Figure˜3. We discuss in more detail this fundamental parametrization of the Ascona model in Section˜3.2.

### v2

@@ Abstract
We introduce a queueing-based sampling framework for continuous-time temporal networks. We focus on a Markovian parametrization in which link start times follow a homogeneous Poisson process and link durations are exponentially distributed. We derive stochastic properties of the resulting link streams and exploit them to generate synthetic temporal networks with controllable smoothness and prescribed event patterns, relevant for the validation and interpretation of methods for community, scale, change-point, and periodicity detection. By coupling this temporal mechanism with block-structured endpoint distributions, we obtain a continuous-time analogue of stochastic block models. We also discuss extensions of the framework, including discrete-time and instantaneous-contact limits.
@@ 1 Introduction
Temporal (or dynamic) networks/graphs are extensions of graphs that incorporate temporal information. Various mathematical objects can be considered to be temporal networks, and different formalisms have been proposed to describe them (e.g., [8, 17, 20]). These objects can be classified on a spectrum of increasing complexity based on the granularity of the encoded temporal information [29, 28]. On one side, we have static graphs, which aggregate in one structure all temporal information; on the other side, there are stream graphs and link streams that avoid any aggregation. Somewhere in between the two sides lie snapshot networks and discrete-time stream graphs/link streams.
@@
Temporal networks are relevant in modeling many complex systems [24], and, as for their static counterparts, understanding their communities/clusters is a fundamental problem in network science. The community structure evolution is typically characterized by archetypal structural changes [10], called events, or combinations thereof [14]. Therefore, sampling methods that allow modeling these archetypes are of paramount importance for the generation of quality synthetic data with controlled ground-truth. One needs these synthetic networks, for instance, for the validation and interpretation of algorithms tasked for community, scale, change-point, and periodicity detection. A review of the main existing methods, particularly relevant for community detection benchmarks, can be found in [9].
@@
For some applications, one would also want to sample networks that are less artificial and instead are closer to real-world datasets. Recently proposed techniques in this direction (e.g., [15, 23, 26]) aim at producing surrogate temporal graphs, structures that maintain some or most properties of a reference (often real-world) temporal network. However, these methods do not provide sufficient control to design a controlled experimental setup, as they bypass the explicit formulation of a modeling mechanism by relying on a pre-existing structure.
@@
Agent-based models (e.g., [25, 32]) are an alternative option. They leverage control of microscopic dynamics of a complex system and combine them so that, in the aggregate, complex phenomena emerge. However, often, the understanding of these models relies heavily on simulations, as the underlying dynamics are complex, leading to mathematical intractability, hence limited stochastic and statistical control.
@@
Most of the temporal network literature, especially on sampling methods, has focused on snapshot networks or discrete-time structures, leaving out the more complex continuous-time ones (a notable exception being [6]). The main contribution of this article is the introduction of a sampling method for continuous-time temporal networks, and the explanation of how to use it to sample data reflecting the above-mentioned event patterns, plus a continuous-time generalization of the stochastic block model (SBM) [1].
@@
The model works in two fundamental steps: in the first, we sample when nodes are connected with a link, and in the second, we sample which nodes to assign to each link sampled in the previous step. Separating the temporal from the connectivity aspect offers benefits in terms of implementation and clarity, while not sacrificing expressivity. This separation mechanism also helps to structure the mathematical analysis of a temporal network (model), as the two aspects can be considered orthogonal to each other, hence treated separately, as in [4]. Connectivity can exhibit almost arbitrary levels of complexity and is responsible for various structural aspects such as assortativity, disassortativity, and nestedness. A complete analysis of its properties from a stochastic and statistical perspective is therefore challenging. The temporal aspect appears to be simpler; its probabilistic and statistical analysis can be studied [4, 30, 31]. This article follows the footsteps of these works: its key idea is to describe the appearance and disappearance of nodes’ connections in terms of queueing processes, and then focus on the evolution of the total number of active connections as a function of time to derive probabilistic and statistical results.
@@
Queueing processes were originally introduced to describe the flow of customers (jobs, particles, calls) arriving at a service facility. It is possible to approximate the evolution of most queueing systems with Markov chains. The advantage of these is that it is possible to calculate the stationary distribution of a chain by solving a linear system of equations called the balance conditions [34].
@@
The structure of this work is as follows. In Section˜2 we introduce the notation and describe the objects studied in the article. It is divided into two parts: the first is devoted to formalisms of temporal networks, the second to queueing processes. In Section˜3 we introduce the Ascona model, a temporal networks model based on queueing theory, and then use a Markovian parametrization of it to generate link streams with desired temporal evolutions. Based on this Markovian framework, we describe a link stream generalization of SBMs, and compare it to dynamic SBMs, a popular temporal extension of them. We conclude the section with a description of variations of the Ascona model. The end of the article, Section˜4, contains a discussion of the presented results and highlights possible future research directions stemming from this work.
@@ 2 Theoretical Framework
In this section, we describe the notation and objects we need to introduce the Ascona model. In the first part, we describe various formalisms of temporal networks following the framework presented in [20], and in the second part, general queueing processes, focusing on the $M/M/\infty$ queue, also known as the ample-server model [16, Ch. 2.7].
@@ 2.1.1 Link Streams
We say that a stream $L^{\prime}=(T^{\prime},V^{\prime},E^{\prime})$ is a sub-stream of $L=(T,V,E)$ if, intuitively, $T^{\prime}\subset T,V^{\prime}\subset V,E^{\prime}\subset E$.
@@ 2.1.2 Snapshot Networks
The length of each slice is determined based on a compromise: one needs slices large enough to ensure that each snapshot captures significant structural information, but large slices lead to losses of temporal information, since all interactions within the same slice are merged. In addition, several or even varying slice durations may be relevant [20].
@@ 2.2 The $M/M/\infty$ Queue
Kendall’s notation, originating from [18], concisely encodes these features as
@@
We are mainly interested in one specific queue, the $M/M/\infty$ one. The ‘$M$’ is a standard symbol in Kendall notation indicating Markovian processes; therefore, the model can be interpreted as a birth (or immigration)-death process where arrivals form a Poisson process with rate $\lambda>0$, and deaths are i.i.d. exponential with rate $\mu>0$ [3]. There are countably infinitely many servers, hence no waiting occurs: each arrival immediately starts the death process.
@@
$p_{m}^{\prime}(t)$ $=\lambda p_{m-1}(t)-(\lambda+m\mu)p_{m}(t)+(m+1)\mu p_{m+1}(t),\qquad m\geqslant 1.$ (2.11)
@@ 2.2.1 Process Evolution
It is known (e.g., [27, Ch. 6]) that, assuming that $M(0)=0$, $M(t)$ is Poisson with a time–dependent mean
@@
More generally, if we start the process at a starting time $t_{s}$, and stop sampling elements after a certain time $t_{e}$, then the process is shifted and truncated, and the mean follows the equation
$m(t)=\begin{cases}0,&t<t_{s},\\[4.0pt] \frac{\lambda}{\mu}\left(1-e^{-\mu(t-t_{s})}\right),&t_{s}\leq t\leq t_{e},\\[4.0pt] \frac{\lambda}{\mu}\left(e^{-\mu(t-t_{e})}-e^{-\mu(t-t_{s})}\right),&t>t_{e}.\end{cases}$ (2.18)
@@ 2.2.2 Stationary Regime
One could also recover $\mathbb{E}[M(\infty)]=\frac{\lambda}{\mu}$ directly from Little’s law (see [22] and [34, Thm. 3.4.1]).
We now derive an estimate for the time $t^{\star}$ at which we can assume that $m(t)$ is close enough to the stationary value. We use this estimate later in Section˜3.2.3 and Section˜3.3.
@@ 2.2.3 Head of the Queue
We consider an $M/M/\infty$ queue that starts empty at time $t=0$, with arrival rate $\lambda>0$ (Poisson arrivals) and i.i.d. service times $S\sim\mathrm{Exp}(\mu)$ of mean $\frac{1}{\mu}>0$.
@@
In the case of $N$ queue samples, the standard deviation of the sample mean is $\sqrt{\frac{m(t)}{N}}$, so it takes more time to establish closeness to stationarity. The computation is equivalent to that of the single sample case. This heuristic argument can be made rigorous in terms of the total variation distance, but a more refined estimation based on mixing times [21] is possible, allowing one to control closeness to stationarity to an arbitrary level of precision.
@@ 2.2.4 Number of Elements in the Queue
The number of elements that join the queue in a time interval $t_{1},t_{2}]$ is, by design, Poisson distributed with average $\lambda(t_{2}-t_{1})$. The number of elements in the queue in the same interval is slightly different. In addition to new elements joining in that interval, we need to count how many started before and ended after $t_{1}$. This amounts approximately to the sum of two independent random variables, one Poisson with average $\lambda(t_{2}-t_{1})$ and the other Poisson with average $m(t_{1})$, which by superposition is again Poisson with average $\lambda(t_{2}-t_{1})+m(t_{1})$. In the stationary regime, this becomes a Poisson random variable with average $\lambda(t_{2}-t_{1})+\lambda/\mu$. We use these facts later in Section˜3.3.
3 The Ascona Model and EDLDE Structures
This section introduces a link stream model and a practical sampling procedure. In the absence of a previously established name, we refer to it as the Ascona model for clarity and convenience. To avoid ambiguity, we explicitly distinguish the model (a probability distribution on link streams) from the sampler (an algorithm that draws approximately from that distribution).
@@ 3.1 The Ascona Model
Let $V=\{v_{1},\dots,v_{N}\}$ be a vertex set and $T\subset\mathbb{R}$ a time domain. Following Section˜2.1.1, a (simple) link stream is a triplet
$L=(T,V,E),\qquad E\subseteq\bigl\{([\alpha,\omega),\{u,v\}):\alpha\leq\omega,\ u\neq v\bigr\},$
so that $([\alpha,\omega),\{u,v\})\in E$ means that the link $\{u,v\}$ is present for all $t\in[\alpha,\omega)$. We denote by
$M_{L}(t)\;\coloneqq\;\bigl|\{\{u,v\}:([\alpha,\omega),\{u,v\})\in E,\ t\in[\alpha,\omega)\}\bigr|$ (3.1)
the number of active links at time $t$.
A convenient way to specify and sample $E$ is via a marked point process of the link start times. Let $\{T_{k}\}_{k\geq 1}$ be random start times on $T$, and let marks be $(D_{k},\{U_{k},V_{k}\})$ with $D_{k}>0$ and $U_{k}\neq V_{k}$. Each $T_{k}$ generates one element
$([T_{k},T_{k}+D_{k}),\{U_{k},V_{k}\})\in E.$
This representation is equivalent to the interval-based definition above, but is more convenient for the Ascona model construction.
The model generates a random link stream by combining:
  1. 1.
    a temporal mechanism producing a (multi)set of link intervals, and
  2. 2.
    a connectivity mechanism assigning an unordered pair of nodes to each interval.
The Ascona model assumes that when and for how long links occur is decoupled from which node pairs interact. Formally, the endpoint marks are independent of the temporal marks:
$\{\{U_{k},V_{k}\}\}_{k\geq 1}\ \perp\!\!\!\perp\ \{(T_{k},D_{k})\}_{k\geq 1}.$ (3.2)
Equivalently, conditional on the realized collection of intervals $\mathcal{I}=\{[T_{k},T_{k}+D_{k})\}_{k\geq 1}$, the unordered pairs $\{U_{k},V_{k}\}_{k\geq 1}$ are i.i.d. draws from a fixed distribution on $\binom{V}{2}$ that does not depend on $\mathcal{I}$. We now describe both mechanisms in more detail.
3.1.1 Temporal Mechanism
In the Ascona model, the temporal structure is defined by a queueing process $A/S/c/K/\bar{K}/D$ on a base time interval $0,t_{e}]$, followed by its tail, obtained by stopping admissions after $t_{e}$ while letting in-service customers complete. A realized queue trajectory produces a collection of active intervals $\mathcal{I}=\{[T_{k},E_{k})\}$ (“customers”) and therefore an occupancy process $M(t)$ (number of simultaneously active customers). Here $T_{k},E_{k})$ corresponds to the generic interval notation $\alpha,\omega)$ used in Section˜2.1.1.
In practice, one often selects only a portion of the queue trajectory (e.g. head, tail, or an approximately stationary segment), as described later in Section˜3.2.3. Additionally, a block can be placed anywhere in the global time domain by applying a time shift $\tau_{a}$ (defined in eq. (3.5)). In this work, we treat trimming and placement as user-chosen design decisions; one could alternatively randomize them.
3.1.2 Connectivity Mechanism
By the Ascona decoupling assumption (eq. (3.2)), after sampling link intervals, endpoints are assigned stochastically and independently across events, according to an endpoint law that is independent of the temporal realization. We consider two equivalent parametrizations of the endpoint law on unordered pairs $\binom{V}{2}$.
Connectivity probability matrix (one-step)
Let $\pi$ be a probability distribution on unordered pairs $\binom{V}{2}$. We encode $\pi$ by a symmetric matrix $P\in[0,1]^{N\times N}$ with $P_{ii}=0$ and
$P_{ij}=P_{ji},\qquad\sum_{1\leq i<j\leq N}P_{ij}=1.$ (3.3)
In this representation, for $i<j$ we have $\pi(\{v_{i},v_{j}\})=P_{ij}$. We call $P$ the connectivity probability matrix.
Connectivity conditional probability matrix (two-step)
Alternatively, sample a first endpoint $U\sim p$ for a distribution $p$ on $V$, then sample a second endpoint $V\sim Q_{U,\cdot}$ where $Q$ is row-stochastic with $Q_{ii}=0$. This factorization is convenient to separate heterogeneous activity ($p$) [25] from mixing patterns ($Q$).
The pair $(p,Q)$ induces a distribution on unordered pairs by symmetrization:
$\pi(\{v_{i},v_{j}\})\;=\;\frac{1}{2}\Big(p_{i}\,Q_{ij}+p_{j}\,Q_{ji}\Big).$ (3.4)
The one-step matrix $P$ corresponding to $(p,Q)$ is given by $P_{ii}=0$ and, for $i<j$, $P_{ij}=\pi(\{v_{i},v_{j}\})$ (hence $P_{ij}=P_{ji}$).
A convenient choice for $p$ is the uniform distribution, which has maximal entropy and therefore can be interpreted as uninformative from a Bayesian perspective.
The practical advantage of the Ascona framework is that complex, nonstationary activity profiles can be built by composing simpler sub-stream units called queue blocks. This makes model design modular: temporal structure is engineered at the block level and then combined together.
3.1.3 Ascona Sampler: Queue Blocks
A queue block $B$ is a link stream sampled on a base interval $0,t_{e}]$ in two stages:
  1. 1.
    Temporal stage: sample a collection of intervals $\mathcal{I}_{B}=\{[T_{k},E_{k})\}_{k=1}^{n_{B}}$ from the chosen queue model, optionally trimmed (head/tail/stationary selection).
  2. 2.
    Connectivity stage: for each $T_{k},E_{k})\in\mathcal{I}_{B}$, sample an unordered pair $\{U_{k},V_{k}\}$ i.i.d. from the chosen endpoint law (either $\pi$ or $(p,Q)$), independently of $\mathcal{I}_{B}$.
For a shift $a\in\mathbb{R}$ define the time shift operator
$\tau_{a}(L)\coloneqq(T+a,V,E_{a}),\qquad E_{a}\coloneqq\{([\alpha+a,\omega+a),\{u,v\}):([\alpha,\omega),\{u,v\})\in E\}.$ (3.5)
Given two streams $L_{1}=(T_{1},V,E_{1})$ and $L_{2}=(T_{2},V,E_{2})$ on the same node set, define their superposition
$L_{1}\oplus L_{2}\;\coloneqq\;(T_{1}\cup T_{2},\ V,\ E_{1}\cup E_{2}),$ (3.6)
interpreting $E_{1}\cup E_{2}$ either as a set (simple stream) or as a multiset (allowing duplicates), depending on the application. Concatenation is the special case where the time supports are disjoint.
If blocks $B_{1},\dots,B_{R}$ are placed at times $a_{1},\dots,a_{R}$, the global stream is
$L\;=\;\bigoplus_{r=1}^{R}\tau_{a_{r}}(B_{r}).$ (3.7)